A particle is moving along the x-axis. Its position as a function of time is given as

π‘₯ = 𝑏𝑑 βˆ’ 𝑐𝑑2


.

a) What must be the units of the constants b and c, if x is in meters and t in seconds?

b) At time zero the particle is at the origin. At what later time t does it pass the origin again?

c) Derive an expression for the x- component of velocity.

d) At what time t is the particle momentarily at rest?

e) Derive an expression for the x-component of the particle’s acceleration, ax.


1
Expert's answer
2021-12-09T15:48:13-0500

Given:


"\ud835\udc65 = \ud835\udc4f\ud835\udc61 \u2212 \ud835\udc50\ud835\udc61^2"

a)

"[b]={\rm m/s},\quad [c]={\rm m/s^2}"

b)

"\ud835\udc65 = \ud835\udc4f\ud835\udc61 \u2212 \ud835\udc50\ud835\udc61^2=0"

"t_1=0,\quad t_2=b/c"

c)

"v_x=x'=b-2ct"

d)

"b-2ct=0"

"t_3=b\/(2c)"

e)

"a_x=v_x'=-2c"